### 04 Mar 2024 v4

# Linking numbers of modular knots

Abstract.
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ acts on the hyperbolic plane $\mathbb{H}\mathbb{P}$ with quotient the modular surface $\mathbb{M}$, whose unit tangent bundle $\mathbb{U}$ is a $3$-manifold homeomorphic to the complement of the trefoil knot in the sphere $\mathbb{S}^{3}$. The hyperbolic conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the closed oriented geodesics in $\mathbb{M}$. Those lift to the periodic orbits for the geodesic flow in $\mathbb{U}$, which define the modular knots.
The linking numbers between modular knots and the trefoil are well understood. Indeed, É. Ghys showed in 2006 that they are given by the Rademacher invariant of the corresponding conjugacy classes. The Rademacher function $\operatorname{Rad}\colon\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}$ is a homogeneous quasi-morphism which he recognised with J. Barge in 1992 as half the primitive of the bounded Euler class.
We are concerned with the linking numbers between modular knots and derive several formulae with arithmetical, combinatorial, topological and group theoretical flavours. In particular we associate to a pair of modular knots a function defined on the character variety of $\operatorname{PSL}_{2}(\mathbb{Z})$, whose limit at the boundary point recovers their linking number. Moreover, we show that the linking number with a modular knot minus that with its inverse yields a homogeneous quasi-morphism on the modular group, and how to extract a free basis out of these.
Acknowlegements
This paper contains the main results obtained in the second part of my thesis. I would thus like to thank my thesis advisors Etienne Ghys and Patrick Popescu-Pampu for their guidance and encouragement; as well as Francis Bonahon, Louis Funar, Jean-Pierre Otal and Anne Pichon who refereed and carefully read my work. I am also grateful to Pierre Dehornoy for sharing his knowledge of modular knots, as well as to the members of the arithmetics and dynamics teams at PSU for inviting me to present my works in their seminars. I owe Marie Dossin for helping me with the figures in tikz. Finally, I thank the referee for their relevant suggestions and questions.
0. Introduction
Context and motivation
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ acts properly discontinuously on the hyperbolic plane $\mathbb{H}\mathbb{P}$ with quotient the modular orbifold $\mathbb{M}$, a hyperbolic surface with conical singularities $i$ and $j$ of order $2$ and $3$, and a cusp $\infty$. The free homotopy classes of loops in $\mathbb{M}$ correspond to the conjugacy classes of its fundamental group $\pi_{1}(\mathbb{M})=\operatorname{PSL}_{2}(\mathbb{Z})$. In particular the hyperbolic conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the closed oriented geodesics in $\mathbb{M}$, called modular geodesics. For hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ the modular geodesic $\alpha$ has length $\lambda_{A}$ equal to the logarithm of the ratio between its eigenvalues $\epsilon_{A}^{\pm 1}$, in formula:
$\operatorname{disc}(A)=\left(\epsilon_{A}-\epsilon_{A}^{-1}\right)^{2}=(\operatorname{Tr}A)^{2}-4=\left(2\sinh\tfrac{1}{2}\lambda_{A}\right)^{2}$
We denote by $I(A,B)$ the geometric intersection number between the associated modular geodesics.
The unit tangent bundle $\mathbb{U}=\operatorname{PSL}_{2}(\mathbb{Z})\backslash\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{M}=\operatorname{PSL}_{2}(\mathbb{Z})\backslash\mathbb{H}\mathbb{P}$ is a $3$-manifold, and the closed oriented geodesics in $\mathbb{M}$ lift to the periodic orbits for the geodesic flow in $\mathbb{U}$. Hence the primitive hyperbolic conjugacy classes in $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the so called modular knots in $\mathbb{U}$, which form the components of the master modular link. The structure of the Seifert fibration $\mathbb{U}\to\mathbb{M}$ reveals that $\mathbb{U}$ is homeomorphic to the complement of a trefoil knot in the sphere. In particular, one may speak of the linking numbers between modular knots and the trefoil and between one another.
Let us recall a combinatorial parametrization of the infinite order conjugacy classes in $\operatorname{PSL}_{2}(\mathbb{Z})$. The Euclidean algorithm shows that the group $\operatorname{SL}_{2}(\mathbb{Z})$ is generated by the transvections
$L=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\qquad\mathrm{and}\qquad R=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$
and more precisely that its submonoid $\operatorname{SL}_{2}(\mathbb{N})$ of matrices with non-negative entries is freely generated by $\{L,R\}$. This submonoid can be identified with its image $\operatorname{PSL}_{2}(\mathbb{N})\subset\operatorname{PSL}_{2}(\mathbb{Z})$. In $\operatorname{PSL}_{2}(\mathbb{Z})$, the conjugacy class of an infinite order element intersects $\operatorname{PSL}_{2}(\mathbb{N})$ along all cyclic permutations of a non-empty $\{L,R\}$-word. The conjugacy class is primitive if and only if the cyclic word is primitive, and it is hyperbolic when the cyclic word contains both letters $L$ and $R$.
One may try to relate the geometry and topology of the master modular link with the arithmetics and combinatorics of conjugacy classes in the modular group. Our previous work [Sim23] relates the geometry of modular geodesics (angles of intersection and lengths of ortho-geodesics) in terms of the arithmetics of conjugacy classes in the modular group (discriminants, cross-ratios between fixed points on the projective line, and their Hilbert symbols). The main results in this paper will relate the linking numbers of modular knots to the combinatorics of the corresponding cyclic words.
The most immediate measures for the complexity of a binary word are given by the sum and difference between the numbers of letters of each sort. For an infinite order $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ we denote by $\operatorname{len}([A])=\#R+\#L$ the combinatorial length and call $\operatorname{Rad}([A])=\#R-\#L$ the Rademacher number of its conjugacy class. In his paper [Ati87] on the Logarithm of the Dedekind eta function, M. Atiyah identified the Rademacher function with no less than six other important functions appearing in diverse areas of mathematics, showing how omnipresent it is.
The function $\operatorname{Rad}\colon\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}$ is a quasi-morphism, meaning that it has a bounded derivative
$d\operatorname{Rad}\colon\operatorname{PSL}_{2}(\mathbb{Z})\times\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}\qquad d\operatorname{Rad}(A,B)=\operatorname{Rad}(B)-\operatorname{Rad}(AB)+\operatorname{Rad}(A)$
and is homogeneous, meaning that $\operatorname{Rad}(A^{n})=n\operatorname{Rad}(A)$ for infinite order $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ and $n\in\mathbb{Z}$. This enabled É. Ghys and J. Barge to recognise it in [BG92] as half the primitive of the bounded Euler class in $H^{2}_{b}(\operatorname{PSL}_{2}(\mathbb{Z});\mathbb{R})$ and explain its ubiquity. In [Ghy07], É. Ghys showed that the linking number of a modular knot with the trefoil equals its Rademacher invariant, and concluded by asking for arithmetical and combinatorial interpretations of the linking pairing between modular knots.
In this work, we will derive several formulae for those linking numbers, providing bridges between the arithmetics and geometry, the combinatorics and algebra, or the dynamics and topology of the modular group.
The arithmetic and geometry of the cosines
The journey from arithmetics began with our previous work [Sim23]. In particular for a field $\mathbb{K}\supset\mathbb{Q}$, we described when two hyperbolic elements $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ of the same discriminant $\Delta$ are conjugate in $\operatorname{PSL}_{2}(\mathbb{K})$. The obstruction is measured in terms of any intersection angle $\theta$ between their modular geodesics $[\alpha],[\beta]$ by the class of $\left(\cos\tfrac{\theta}{2}\right)^{2}\in\mathbb{K}^{\times}$ modulo the group of norms over $\mathbb{K}$ of the extension $\mathbb{K}[\sqrt{\Delta}]$. When this obstruction vanishes, the elements of $\operatorname{PSL}_{2}(\mathbb{K})$ which conjugate $A$ to $B$ are paremetrized by the points $(X,Y)\in\mathbb{K}$ of the generalised Pell-Fermat conic with equation $X^{2}-\Delta Y^{2}=\left(\cos\tfrac{\theta}{2}\right)^{2}$. Hence the geometric quantities $\left(\cos\tfrac{\theta}{2}\right)^{2}=\tfrac{1+\cos(\theta)}{2}$ given for some representatives $A,B\in\operatorname{SL}_{2}(\mathbb{R})$ whose axes intersect in $\mathbb{H}\mathbb{P}$ by
$\cos(\theta)=\operatorname{sign}(\operatorname{Tr}(A)\operatorname{Tr}(B))\tfrac{\operatorname{Tr}(AB)-\operatorname{Tr}(AB^{-1})}{\sqrt{\operatorname{disc}A\operatorname{disc}B}}$
have an arithmetic meaning, and they will reappear under various forms in the sequel.
Linking functions on the character variety
Let us introduce, for any pair of modular geodesics $[\alpha],[\beta]\subset\mathbb{M}$, the following summations over their oriented intersection angles $\theta\in\,]-\pi,\pi[$:
$\operatorname{L}_{1}(A,B)=\tfrac{1}{2}\sum\left(\cos\tfrac{\theta}{2}\right)^{2}\qquad\mathrm{and}\qquad\operatorname{C}_{1}(A,B)=\tfrac{1}{2}\sum\left(\cos\theta\right)$
and study their variations as we deform the metric on $\mathbb{M}$ by opening the cusp.
The complete hyperbolic metrics on the orbifold $\mathbb{M}$ correspond to the faithful and discrete representations $\rho\colon\operatorname{PSL}_{2}(\mathbb{Z})\to\operatorname{PSL}_{2}(\mathbb{R})$ up to conjugacy. They form a $1$-dimensional real algebraic set parametrized by $q\in\mathbb{R}^{*}$, deforming $L$ and $R$ according to:
$L_{q}=\begin{pmatrix}q&0\\ 1&q^{-1}\end{pmatrix}\qquad\mathrm{and}\qquad R_{q}=\begin{pmatrix}q&1\\ 0&q^{-1}\end{pmatrix}.$
The primitive hyperbolic conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$ still index the hyperbolic geodesics in the quotient $\mathbb{M}_{q}=\rho_{q}(\operatorname{PSL}_{2}(\mathbb{Z}))\backslash\mathbb{H}\mathbb{P}$ which do not surround the cusp. The intersection points between $q$-modular geodesics $[\alpha_{q}],[\beta_{q}]\subset\mathbb{M}_{q}$ persist by deformation, so we may define the analogous sums $\operatorname{L}_{q}(A,B)$ and $\operatorname{C}_{q}(A,B)$ over their intersection angles $\theta_{q}\in\,]-\pi,\pi[$ of the $\tfrac{1}{2}\left(\cos\tfrac{1}{2}\theta_{q}\right)^{2}$ and $\tfrac{1}{2}\left(\cos\theta_{q}\right)$.
As $q\to\infty$, the hyperbolic orbifold $\mathbb{M}_{q}$ has a convex core which retracts onto a thin neighbourhood of the long geodesic arc connecting its conical singularities, whose preimage in the universal cover $\mathbb{H}\mathbb{P}$ is a trivalent tree. In the limit we recover the action of $\operatorname{PSL}_{2}(\mathbb{Z})$ on its Bruhat-Tits building, the infinite planar trivalent tree $\mathcal{T}$, and by studying its combinatorics we shall prove the following.
Theorem 0.1 (Linking and intersection from boundary evaluations).
For primitive hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$, the limits of the functions $\operatorname{L}_{q}(A,B)$ and $\operatorname{C}_{q}(A,B)$ at the boundary point of the $\operatorname{PSL}_{2}(\mathbb{R})$-character variety of $\operatorname{PSL}_{2}(\mathbb{Z})$ recover their linking and intersection numbers:
$\operatorname{L}_{q}(A,B)\xrightarrow[q\to\infty]{}\operatorname{lk}(A,B)$
$\operatorname{C}_{q}(A,B)\xrightarrow[q\to\infty]{}\operatorname{lk}(A,B)-\operatorname{lk}(A^{-1},B)=2\operatorname{lk}(A,B)-\tfrac{1}{2}I(A,B)$
Hence the functions $\operatorname{L}_{q}$ and $\operatorname{C}_{q}$ interpolate between the geometry at $q=1$ of the arithmetic group $\operatorname{PSL}_{2}(\mathbb{Z})\subset\operatorname{PSL}_{2}(\mathbb{R})$ and the topology at $q=\infty$ of the combinatorial action $\operatorname{PSL}_{2}(\mathbb{Z})\to\operatorname{Aut}(\mathcal{T})$.
Linking functions and Alexander polynomials
The graphs of $q\in\mathbb{C}\mapsto\operatorname{L}_{q}(A,B)\in\mathbb{C}$ for various pairs $A,B$ suggest that their zeros tend to accumulate on the unit circle. This reminds us of the various results and conjectures concerning the roots of Alexander polynomials, so we propose a possible thread to follow in this direction.
A primitive hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})=\pi_{1}(\mathbb{M})$ corresponds to a primitive modular geodesic in $\alpha\subset\mathbb{M}$. It lifts to a modular knot in $\vec{\alpha}\subset\mathbb{U}$ which in turns yields a conjugacy class in $\mathcal{B}_{3}=\pi_{1}(\mathbb{U})$. A conjugacy class in the braid group on three strands defines, by taking its closure, a link $\sigma_{A}$ in a solid torus. In [Sim22, Proposition 5.16] we relate the Alexander polynomial $\Delta(\sigma_{A})\in\mathbb{Z}[t^{\pm 1}]$ of this link $\sigma_{A}$ to the Fricke polynomial $\operatorname{Tr}A_{q}\in\mathbb{Z}[q^{\pm 1}]$ of the modular geodesic $[\alpha]$.
Proposition 0.2 (Alexander, Fricke and Rademacher).
For a primitive hyperbolic $A\in\operatorname{SL}_{2}(\mathbb{N})$, the Alexander polynomial of the link $\sigma_{A}$ is given in terms of $q=\sqrt{-t}$ by:
$\Delta(\sigma_{A})=\tfrac{q^{\operatorname{Rad}(A)}-\operatorname{Tr}(A_{q})+q^{-\operatorname{Rad}(A)}}{(q-q^{-1})^{2}}$
Now recall that $\operatorname{C}_{q}(A,B)=\operatorname{L}_{q}(A,B)-\operatorname{L}_{q}(A,B^{-1})$ can be expressed as finite sum of terms:
$\cos(A_{q},B_{q})=\operatorname{sign}(\operatorname{Tr}(A_{q})\operatorname{Tr}(B_{q}))\tfrac{\operatorname{Tr}(A_{q}B_{q})-\operatorname{Tr}(A_{q}B_{q}^{-1})}{\sqrt{\operatorname{disc}A_{q}\operatorname{disc}B_{q}}}\qquad\mathrm{where}\qquad\operatorname{disc}(C_{q})=(\operatorname{Tr}C_{q})^{2}-4$
This is how one may compare the concentration property for the zeros of $\operatorname{L}_{q}$ around the unit circle with those of Alexander polynomials, but we will not pursue this direction any further.
Linking numbers and homogeneous quasi-morphisms
The limiting values $\operatorname{C}_{q}(A,B)\to\operatorname{lk}(A,B)-\operatorname{lk}(A^{-1},B)$ will now provide a bridge from the representation theory to the bounded cohomology of $\operatorname{PSL}_{2}(\mathbb{Z})$. For every group $\Pi$, the real vector space $PX(\Pi)$ of homogeneous quasi-morphisms is a Banach space for the norm $\lVert df\rVert_{\infty}$, as was shown in [MM85, Iva88].
Theorem 0.3.
For every hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$, the function $\operatorname{C}_{A}\colon B\mapsto\operatorname{lk}(A,B)-\operatorname{lk}(A^{-1},B)$ is a homogeneous quasi-morphism $\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}$.
Let $\mathcal{P}$ denote the set of primitive infinite order conjugacy classes in $\operatorname{PSL}_{2}(\mathbb{Z})$, and $\mathcal{P}_{0}$ the subset of those which are stable under inversion. Choose a partition $\mathcal{P}\setminus\mathcal{P}_{0}=\mathcal{P}_{-}\sqcup\mathcal{P}_{+}$ in two subsets in bijection by the inversion. We may choose $R\in\mathcal{P}_{+}$, and denote $\operatorname{C}_{R}:=\operatorname{Rad}$ by convention.
Theorem 0.4.
The collection of $\operatorname{C}_{A}\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ for $A\in\mathcal{P}_{+}$ is linearly independent and every element $f\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ can be written as $f=\sum_{A\in\mathcal{P}_{+}}c_{A}(f).\operatorname{C}_{A}$ for unique $c_{A}(f)\in\mathbb{R}$.
To prove the non-triviality and linear independence of the $\operatorname{C}_{A}$ for $A\in\mathcal{P}_{+}$, we were led to show the non-degeneracy of the linking form, which is interesting in its own right.
Theorem 0.5.
If hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ are link equivalent, namely $\operatorname{lk}(A,X)=\operatorname{lk}(B,X)$ for all hyperbolic $X\in\operatorname{PSL}_{2}(\mathbb{Z})$, then they are conjugate.
The results in this section may be compared to the classical representation theory of compact groups, in which the characters of irreducible representations provide an orthonormal basis for the class functions. Indeed, we have found a family of cosign functions $\operatorname{C}_{A}$ whose periods correspond to the primitive conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$, and they form a basis for the space of quasi-characters $PX(\operatorname{PSL}_{2}(\mathbb{Z}))$, which is in some sense orthogonal with respect to the linking form (but we refer to the proofs in section 7 for a better explanation of this orthogonality).
1. The group $\operatorname{PSL}_{2}(\mathbb{R})$: discriminant and cross-ratio
The group $\operatorname{PSL}_{2}$ acts by conjugacy over itself. In this section, we recall from [Sim23] the main invariants which enable to describe the orbits of single elements and of pairs of elements. Those are the discriminant $\operatorname{disc}(A)$ and the cross-ratio $\operatorname{bir}(A,B)$.
1.1. Over a field of characteristic $\neq 2$
Let $\mathbb{K}$ be a field of characteristic different from $2$. The automorphism group $\operatorname{PGL}_{2}(\mathbb{K})$ of the projective line $\mathbb{K}\mathbb{P}^{1}$ acts freely transitively on triples of distinct points, and the unique algebraic invariant of four points $u,v,x,y\in\mathbb{K}\mathbb{P}^{1}$ is the cross-ratio:
(1) $\operatorname{bir}(u,v,x,y)=\frac{(v-u)}{(v-x)}\div\frac{(y-u)}{(y-x)}\in\mathbb{K}\mathbb{P}^{1}$
It satisfies in particular $\operatorname{bir}(z,0,1,\infty)=z$ and $\operatorname{bir}(z,0,w,\infty)=z/w$, whence the cocycle rule:
$\operatorname{bir}(z,v,x,y)=\operatorname{bir}(z,v,w,y)\operatorname{bir}(w,v,x,y).$
For a triple $(u,v,w)$ of points in $\mathbb{K}\mathbb{P}^{1}$, we define their Maslov index in $\{0\}\sqcup\mathbb{K}^{\times}/(\mathbb{K}^{\times})^{2}$ by lifting them to vectors $(\vec{u},\vec{v},\vec{w})$ in $\mathbb{K}^{2}$ which are not all zero but with zero sum, and taking the determinant $\det(\vec{u},\vec{v})=\det(\vec{v},\vec{w})=\det(\vec{w},\vec{u})$. It is preserved by the subgroup $\operatorname{PSL}_{2}(\mathbb{K})$, which acts freely transitively on the triples of distinct points with a given Maslov index.
A non-trivial $A\in\operatorname{PSL}_{2}(\mathbb{K})$ has two fixed points $\alpha_{-},\alpha_{+}$ in the projective line over $\sqrt{\mathbb{K}}$, that are well defined up to permutation. The element $\epsilon_{A}^{\pm 2}:=\operatorname{bir}(Ax,\alpha_{\mp},x,\alpha_{\pm})$ does not depend on $x\in\mathbb{K}\mathbb{P}^{1}\setminus\{\alpha_{-},\alpha_{+}\}$, and is well defined up to inversion: it is called the period of $A$. The unique algebraic function on $\operatorname{PSL}_{2}(\mathbb{K})$ which is invariant by conjugacy is the discriminant:
$\operatorname{disc}(A)=\left(\epsilon_{A}-\epsilon_{A}^{-1}\right)^{2}$
whose class in $\{0\}\sqcup\mathbb{K}^{\times}/(\mathbb{K}^{\times})^{2}$ defines the type of $A$. All non-trivial elements with $\operatorname{disc}=0$ are conjugate. The elements with $\operatorname{disc}\neq 0$ are called semi-simple. The conjugacy classes of elements with $\operatorname{disc}\equiv 1\bmod{(\mathbb{K}^{\times})^{2}}$ are uniquely characterised by the value of their discriminant.
Consider $A,B\in\operatorname{PSL}_{2}(\mathbb{K})$ of the same type, and fix a square root of $\operatorname{disc}(A)\operatorname{disc}(B)\in(\mathbb{K}^{\times})^{2}$. Then one may order their fixed points $(\alpha_{-},\alpha_{+})$ and $(\beta_{-},\beta_{+})$ up to simultaneous inversion, and consistently define their cross-ratio $\operatorname{bir}(A,B)$ by:
$\operatorname{bir}(A,B):=\operatorname{bir}(\alpha_{+},\alpha_{-},\beta_{+},\beta_{-})\in\mathbb{K}\mathbb{P}^{1}$
which is $\notin\{0,1,\infty\}$ unless $A$ and $B$ share a fixed point, and satisfies the symmetry property:
$\frac{1}{\operatorname{bir}(A,B)}+\frac{1}{\operatorname{bir}(A,B^{-1})}=1.$
We may also define their cosine (using their adjoint action on the Lie algebra $\operatorname{\mathfrak{sl}}_{2}(\mathbb{K})$ as in [Sim23]), which is related to the cross-ratio by:
(2) $\cos(A,B)=\frac{1}{\operatorname{bir}(A,B)}-\frac{1}{\operatorname{bir}(A,B^{-1})}.$
Theorem 1.1 ([Sim23]).
Two semi-simple elements $A,B\in\operatorname{PSL}_{2}(\mathbb{K})$ are conjugate if and only if $\operatorname{disc}(A)=\Delta=\operatorname{disc}(A)$ and $\operatorname{bir}(A,B)\equiv 1\bmod{\operatorname{Norm}_{\mathbb{K}}\mathbb{K}[\sqrt{\Delta}]}$.
More precisely, if $\Delta\neq 0$ and $\operatorname{bir}(A,B)\notin\{1,\infty\}$, then we may lift $A,B$ in $\operatorname{SL}_{2}(\mathbb{K})$ so that $\mathfrak{a}=A-\tfrac{1}{2}\operatorname{Tr}(A),\mathfrak{b}=B-\tfrac{1}{2}\operatorname{Tr}(A)$ satisfy $\operatorname{Tr}(\mathfrak{a}\mathfrak{b})=-\Delta\cos(A,B)$, and the elements $C\in\operatorname{SL}_{2}(\mathbb{K})$ such that $CAC^{-1}=B$ are parametrized by the $\mathbb{K}$-points of the Pell-Fermat conic:
$(x,y)\in\mathbb{K}\times\mathbb{K}\>\colon\>\quad(2x)^{2}+\Delta y^{2}=\operatorname{bir}(A,B)\quad\text{according to}\quad C(x,y)=x(\mathfrak{1}+\mathfrak{b}\mathfrak{a}^{-1})+y(\mathfrak{a}+\mathfrak{b}).$
1.2. Over the real field
The automorphism group $\operatorname{PGL}_{2}(\mathbb{C})\simeq\operatorname{PSL}_{2}(\mathbb{C})$ of the complex projective line $\mathbb{C}\mathbb{P}^{1}$ contains $\operatorname{PGL}_{2}(\mathbb{R})$ as the stabiliser of the real projective line $\mathbb{R}\mathbb{P}^{1}$. The index-two subgroup $\operatorname{PSL}_{2}(\mathbb{R})$ also preserves the upper half-plane =$\mathbb{H}\mathbb{P}=\{z\in\mathbb{C}\mid\Im(z)>0\}\subset\mathbb{C}\mathbb{P}^{1}$, or equivalently the orientation induced on its boundary $\partial\mathbb{H}\mathbb{P}$, also given by the cyclic order $\operatorname{cord}(x,y,z)\in\{\pm 1\}$ of any triple of distinct points $x,y,z$ of $\mathbb{R}\mathbb{P}^{1}$ (which equals their Maslov index).
The complex structure on $\mathbb{H}\mathbb{P}$ is conformal to a unique hyperbolic metric. The hyperbolic distance $\lambda$ between $w,z\in\mathbb{H}\mathbb{P}$ can be deduced from the cross-ratio by $\operatorname{bir}(\bar{z},z,\bar{w},w)^{-1}=\left(\cosh\tfrac{\lambda}{2}\right)^{2}$. This realises $\operatorname{PSL}_{2}(\mathbb{R})$ as the positive isometry group of the hyperbolic plane: it preserves the previous cross-ratio and acts simply-transitively on positive triples of distinct points in $\mathbb{R}\mathbb{P}^{1}$, thus it preserves the hyperbolic metric and acts simply transitively on the unit tangent bundle of $\mathbb{H}\mathbb{P}$.
The type of $A\in\operatorname{PSL}_{2}(\mathbb{R})$ is elliptic or parabolic or hyperbolic according to the value of $\operatorname{sign}\operatorname{disc}(A)\in\{-1,0,1\}$, equal to the number of distinct fixed points in $\mathbb{R}\mathbb{P}^{1}$ minus $1$.
A hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{R})$ acts on $\mathbb{H}\mathbb{P}$ by translation along an (oriented) geodesic $\alpha$ whose endpoints $\alpha_{-},\alpha_{+}\in\mathbb{R}\mathbb{P}^{1}$ are its repulsive and attractive fixed points. With this order the period satisfies $\epsilon_{A}^{2}>1$. The translation length $\lambda_{A}=\log(\epsilon_{A}^{2})$ yields $\operatorname{disc}(A)=\left(2\sinh\tfrac{1}{2}\lambda_{A}\right)^{2}$.
Lemma 1.2 (Real geometry of the cross-ratio).
Consider hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{R})$ whose axes $\alpha,\beta$ have distinct endpoints. For any lifts $A,B\in\operatorname{SL}_{2}(\mathbb{R})$ we have:
$\cos(A,B)=\operatorname{sign}(\operatorname{Tr}(A)\operatorname{Tr}(B))\frac{\operatorname{Tr}(AB)-\operatorname{Tr}(AB^{-1})}{\sqrt{\operatorname{disc}(A)\operatorname{disc}(B)}}.$
If $\alpha$ and $\beta$ intersect, then it is with an angle $\theta\in]-\pi,\pi[$ oriented anti-clockwise from $\alpha$ to $\beta$ according to $\operatorname{sign}(\theta)=:\operatorname{sign}(A,B)$, and satisfying $\cos(\theta)=\cos(A,B)$, thus:
$\tfrac{1}{\operatorname{bir}(A,B)}=\tfrac{1+\cos(\theta)}{2}=\left(\cos\tfrac{\theta}{2}\right)^{2}$
Note that $\theta\mapsto-\theta$ by transposition of $(A,B)$ whereas $\theta\mapsto\operatorname{sign}(\theta)\pi-\theta$ by inversion of $A$ or $B$.
If $\alpha$ and $\beta$ are disjoint, then they are joint by a unique ortho-geodesic arc with length $\lambda\in\mathbb{R}$, whose $\operatorname{sign}(\lambda)=\operatorname{sign}\operatorname{bir}(A,B)$ compares the co-orientations induced by each axis, and satisfying $\cos(A,B)=\operatorname{sign}(\lambda)\cosh(\lambda)$, thus respectively:
$\tfrac{1}{\operatorname{bir}(A,B)}=\tfrac{1+\cosh(\lambda)}{2}=\left(\cosh\tfrac{\lambda}{2}\right)^{2}\quad\mathrm{and}\quad\tfrac{1}{\operatorname{bir}(A,B)}=\tfrac{1-\cosh(\lambda)}{2}=-\left(\sinh\tfrac{\lambda}{2}\right)^{2}$
Note that $\lambda$ is invariant by transposition of $(A,B)$ whereas $\lambda\mapsto-\lambda$ by inversion of $A$ or $B$.
$\alpha_{-}$$\alpha_{+}$$\beta_{-}$$\beta_{+}$$\theta$
$\alpha_{-}$$\beta_{+}$$\beta_{-}$$\alpha_{+}$$\lambda$
$\alpha_{-}$$\beta_{-}$$\beta_{+}$$\alpha_{+}$$\lambda$
Angle $\theta\in\,]0,\pi[$. Ortho-geodesics: co-oriented ($\lambda>0$) and disco-oriented ($\lambda<0$).
Proof.
The statements are invariant by the $\operatorname{PGL}_{2}(\mathbb{R})$ action so we may fix three points among $(\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})$ and compute in terms of the last. The cosine formula was used in [Wol81]. ∎
2. The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$
2.1. The modular orbifold
The subgroup $\operatorname{PSL}_{2}(\mathbb{Q})$ of $\operatorname{PSL}_{2}(\mathbb{R})$ is the stabiliser of the rational projective line $\mathbb{Q}\mathbb{P}^{1}$. The discrete subgroup $\operatorname{PSL}_{2}(\mathbb{Z})$ is the stabiliser of the ideal triangulation $\triangle$ of $\mathbb{H}\mathbb{P}$ with vertex set $\mathbb{Q}\mathbb{P}^{1}$ and edges all geodesics whose endpoints $\tfrac{p}{q},\tfrac{r}{s}$ satisfy $\lvert ps-qr\rvert=1$.
Consider the action of $\operatorname{PSL}_{2}(\mathbb{Z})$ on $\triangle$. It is transitive on the set of edges, which is in bijection with the orbit of $i\in(0,\infty)$, and the stabiliser of $i$ is the subgroup of order $2$ generated by $S$. It is transitive on the set of triangles, which is in bijection with the orbit of $j=\exp(i\pi/3)$, and the stabiliser of $j$ is the subgroup of order $3$ generated by $T$. Thus it is freely transitive on the flags of $\triangle$, or equivalently on the oriented edges, and we deduce that $\operatorname{PSL}_{2}(\mathbb{Z})=\mathbb{Z}/2*\mathbb{Z}/3$ is the free amalgam of its subgroups generated by $S$ and $T$.
$S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\qquad T=\begin{pmatrix}1&-1\\ 1&0\end{pmatrix}$
We also find that $\operatorname{PSL}_{2}(\mathbb{Z})$ acts properly discontinuously on $\mathbb{H}\mathbb{P}$ with fundamental domain the triangle $(\infty,0,j)$. We may cut it along the geodesic arc $(i,j)$ to obtain a pair of isometric triangles $(i,j,\infty)$ and $(i,j,0)$. Identifying them along their isometric edges yields the modular orbifold
$\mathbb{M}=\operatorname{PSL}_{2}(\mathbb{Z})\backslash\mathbb{H}\mathbb{P}.$
It is a hyperbolic two-dimensional orbifold, with conical singularities of order $2$ and $3$ associated to the fixed points $i$ and $j$ of $S$ and $T$, and a cusp associated to the fixed point $\infty\in\partial\mathbb{H}\mathbb{P}$ of $R=TS^{-1}$.
$\frac{3}{1}$$\frac{1}{3}$$-\frac{1}{3}$$-\frac{3}{1}$$\frac{3}{2}$$\frac{2}{3}$$-\frac{3}{2}$$-\frac{2}{3}$$-\frac{2}{1}$$\frac{1}{0}$$\frac{1}{1}$$\frac{0}{1}$$-\frac{1}{1}$$\frac{2}{1}$$-\frac{1}{2}$$\frac{1}{2}$
[…]
The ideal triangulation of $\mathbb{H}\mathbb{P}$ together with its dual trivalent tree $\mathcal{T}$ yield the modular tessellation with fundamental domain $(0,j,\infty)$.
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ is the orbifold fundamental group of $\mathbb{M}$, so its conjugacy classes correspond to the free homotopy classes of (oriented) loops in $\mathbb{M}$. The elliptic conjugacy classes are those of $S$ and $T^{\pm 1}$ which correspond to loops encircling the singularities, and the parabolic conjugacy classes are those of $R^{n}$ which correspond to loops encircling the cusp. The conjugacy class of a hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ corresponds to the free homotopy class of a unique closed geodesic $[\alpha]\subset\mathbb{M}$, and its length equals $\lambda_{A}=\log(\epsilon_{A}^{2})$. These are the so called modular geodesics.
2.2. Acting on a trivalent tree
The preimage of the segment $(i,j)\subset\mathbb{M}$ in $\mathbb{H}\mathbb{P}$ forms a bipartite tree $\mathcal{T}^{\prime}$, the first barycentric subdivision of a trivalent tree $\mathcal{T}$ which is dual to the ideal triangulation. The base edge $(i,j)$ of $\mathcal{T}^{\prime}$ defines the oriented base edge $\vec{e}_{i}$ of $\mathcal{T}$. The action of $\operatorname{PSL}_{2}(\mathbb{Z})$ is freely transitive on the set of edges of $\mathcal{T}^{\prime}$ hence on the set of oriented edges of $\mathcal{T}$. It also preserves the cyclic order on the set of edges incident to each vertex (given by the surface embedding $\mathcal{T}\subset\mathbb{H}\mathbb{P}$). This is equivalent to the cyclic order function $\operatorname{cord}(x,y,z)\in\{-1,0,1\}$ of three points $x,y,z\in\mathcal{T}\cup\partial\mathcal{T}$. Thus $\operatorname{PSL}_{2}(\mathbb{Z})$ is the full automorphism group of cyclically ordered simplicial tree $(\mathcal{T},\operatorname{cord})$.
We may now use this action to find some conjugacy invariants of primitive elements by considering their stable subsets. Recall that an element in $\operatorname{PSL}_{2}(\mathbb{Z})$ is called primitive when it generates a maximal cyclic subgroup. The (primitive) elliptic conjugacy classes correspond to the vertices of $\mathcal{T}^{\prime}$ and the primitive parabolic conjugacy classes correspond to the connected components of $\mathbb{H}\mathbb{P}\setminus\mathcal{T}$.
Let $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ be primitive of infinite order. It acts on $\mathcal{T}$ by translation along an oriented geodesic $\alpha_{\mathcal{T}}$ called its combinatorial axis, with endpoints $\alpha_{-},\alpha_{+}\in\partial\mathcal{T}=\mathbb{R}\mathbb{P}^{1}$. Observe that $\alpha_{\mathcal{T}}$ passes through the oriented base edge $\vec{e}_{i}$ of $\mathcal{T}$ exactly when its endpoints satisfy $\alpha_{-}\leq 0\leq\alpha_{+}$, which is equivalent to saying that $A$ maps the base triangle $(0,1,\infty)$ to a triangle of the form $(\tfrac{b}{d},\tfrac{a+b}{c+d},\tfrac{a}{c})$ with $a,b,c,d\in\mathbb{N}$, in other terms that $A$ belongs to the monoid $\operatorname{PSL}_{2}(\mathbb{N})$ freely generated by $\{L,R\}$. In that case, $\alpha_{\mathcal{T}}$ follows a periodic sequence of left and right turns given by the $\{L,R\}$-factorisation of $A$, or the continued fraction expansion of the periodic number $\alpha_{+}$.
The conjugacy class of $A$ corresponds to the orbit of $\alpha_{\mathcal{T}}$ under the action of $\operatorname{PSL}_{2}(\mathbb{Z})$ on $\mathcal{T}$. Hence the conjugacy classes of non-elliptic elements in $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the cyclic words over the alphabet $\{L,R\}$, and the hyperbolic classes yield the cycles in which both letters appear. The linear representatives of such an $\{L,R\}$-cycle parametrize the intersection of the corresponding conjugacy class with $\operatorname{PSL}_{2}(\mathbb{N})$, whose elements are called its Euclidean representatives. For infinite order $A\in\operatorname{PSL}_{2}(\mathbb{Z})$, the minimum displacement length $d(x,A\cdot x)$ of a vertex $x\in\mathcal{T}$ equals the combinatorial length $\operatorname{len}(A)=\#R+\#L$ of a Euclidean representative.
Since the combinatorial and geometric axes of a hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ have the same endpoints $\alpha_{-},\alpha_{+}\in\mathbb{R}\mathbb{P}^{1}=\partial\mathbb{H}\mathbb{P}=\partial\mathcal{T}$, they intersect the ideal triangulation in the same pattern. However the geometric axis also contains the information of the intersection pattern with its first barycentric subdivision, which is equivalent to the isotopy class of the modular geodesic in $\mathbb{M}$.
The geometric axes in $\mathbb{P}\mathbb{H}$ and their projections in $\mathbb{M}$ of $RL$, $RLL$, $RLLL$.
In [Sim22, Chapter 3], we recover the isotopy class of $[\alpha]\subset\mathbb{M}$ from the $\{L,R\}$-cycle of $A$. We also describe when $[\alpha]$ passes through the singular points $i$ or $j$. Moreover [Sim22, Lemma 2.27] shows that if a (primitive) hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})$ is stable under inversion, then it contains (exactly $4$) symmetric matrices, and those all lie in $\operatorname{PSL}_{2}(\mathbb{N})$ up to inversion.
3. From the geometric cosine to the combinatorial cosign
3.1. The functions $\operatorname{cross}$ and $\operatorname{cosign}$.
We now use the representation $\operatorname{PSL}_{2}(\mathbb{Z})=\operatorname{Aut}(\mathcal{T},\operatorname{cord})$ to find conjugacy invariants for pairs of primitive infinite order elements by comparing the relative positions of their stable subsets, namely their combinatorial axes.
Let us first derive from the cyclic order function of three points, the crossing function of four points $u,v,x,y\in\mathcal{T}\cup\partial\mathcal{T}$ by:
(3) $\operatorname{cross}(u,v,x,y)=\tfrac{1}{2}\left(\operatorname{cord}(u,x,v)-\operatorname{cord}(u,y,v)\right)$
that is the algebraic intersection number of the oriented geodesics $(u,v)$ and $(x,y)$. We denote $\operatorname{\lvert cross\rvert}(u,v,x,y)\in\{0,\tfrac{1}{2},1\}$ the absolute value of $\operatorname{cross}(u,v,x,y)$ which yields the linking number of the cycles $(u,v)$, $(x,y)$ in the cyclically ordered boundary $\partial\mathcal{T}$. One may compare the formula (3) defining $\operatorname{cross}$ with the formula (1) defining $\operatorname{bir}$ noticing that $\operatorname{cord}(x,y,z)=\operatorname{sign}\tfrac{y-z}{y-x}$. In particular, for $u,v,x,y\in\mathbb{R}\mathbb{P}^{1}$ we have
$\operatorname{\lvert cross\rvert}(u,v,x,y)=1\iff\operatorname{bir}(u,v,x,y)>1$
Now consider two oriented bi-infinite geodesics $\alpha_{\mathcal{T}}$ and $\beta_{\mathcal{T}}$ of $\mathcal{T}$. Their intersection is either empty in which case we define $\operatorname{cosign}(\alpha_{\mathcal{T}},\beta_{\mathcal{T}})=0$, or else it consists in a geodesic containing at least one edge along which we may thus compare their orientations by $\operatorname{cosign}(\alpha_{\mathcal{T}},\beta_{\mathcal{T}})\in\{-1,+1\}$.
The functions $\operatorname{cross}$ and $\operatorname{cosign}$ are $\operatorname{PSL}_{2}(\mathbb{Z})$-invariant, symmetric, and inverting the orientation of one argument results in a change of sign.
Configurations of axes: $\operatorname{cross}$ and $\operatorname{cosign}$. Note that $\operatorname{cross}\neq 0\implies\operatorname{cosign}=\pm 1$.
The identity map and some graphs of $\operatorname{L}_{q}(A,B)$ for $q\in\mathbb{C}$ with $\lvert q\rvert<1+\epsilon$.
2 $\operatorname{cosign}$ $\operatorname{cross}$ $\mathbf{+1}$ $\mathbf{0}$ $\mathbf{-1}$
$\mathbf{+1}$
[…]
$\mathbf{-1}$
[…]
For hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ with axes $\alpha_{\mathcal{T}}=(\alpha_{-},\alpha_{+})$ and $\beta_{\mathcal{T}}=(\beta_{-},\beta_{+})$ in $\mathcal{T}$, we denote by $\operatorname{cross}(A,B)=\operatorname{cross}(\alpha_{+},\alpha_{-},\beta_{+},\beta_{-})$ and $\operatorname{cosign}(A,B)=\operatorname{cosign}(\alpha_{\mathcal{T}},\beta_{\mathcal{T}})$. Notice $\operatorname{cosign}(A,B)=1$ if and only if there exists $C\in\operatorname{PSL}_{2}(\mathbb{Z})$ such that $CAC^{-1},CBC^{-1}\in\operatorname{PSL}_{2}(\mathbb{N})$, in which case the set of such $C$ corresponds to the edges in $\alpha_{\mathcal{T}}\cap\beta_{\mathcal{T}}\subset\mathcal{T}$.
The main observation is that the zeros and poles of $\operatorname{L}_{q}(A,B)$ seem to concentrate on the unit circle. This is neither surprising nor obvious, as we explain in the next paragraph.
Configurations of axes: $\operatorname{cross}$ and $\operatorname{cosign}$. Note that $\operatorname{cross}\neq 0\implies\operatorname{cosign}=\pm 1$.
The identity map and some graphs of $\operatorname{L}_{q}(A,B)$ for $q\in\mathbb{C}$ with $\lvert q\rvert<1+\epsilon$.